# Theory of the Spin Seebeck effect induced by crystal-field excitations in Tb3Fe5O12

### v2

@@ IV Spin current generation in TbIG at low temperatures
To address under which conditions the CFE could contribute to generating a spin current, we adopt a simplified Hamiltonian that includes the essential low energy behaviour of a relatively localized crystal field level coupled to a single dispersive mode of opposite chirality.
@@
This could be considered as an extension to the model adopted in Ref. 10 (see supplemental material therein), to include the magnetic field dependence of the CF excitations (see Fig. 5). There are only two sublattices represented in this model, one for the Fe spins $\vec{s}_{a}$ and one for Tb spins $\vec{S}_{c}$. Here the suffices $a$ and $c$ are to distinguish the distinct Fe and Tb sublattices, respectively, as reminiscent from the Wycoff positions in garnet crystals. Beside the Fe-Fe and Tb-Fe exchanges ($J,J^{\prime}$), both spins are subjected to simplified single-ion anisotropy potentials ($D_{a},D_{c}$) to incorporate the axial constraints known from the garnet materials.
@@
where the $\eta$ summation is taken over the $z_{a}=12$ nearest neighbor $a$ sites of one $a$ site, and the $\tilde{\eta}$ summation is of over the $z_{c}=6$ nearest neighbouring $c$ rare-earth sites of one $a$ site. Note that in Eq. (7) the RE-sites interact only with the n.n. Fe-sites and have no interaction within the same sublattice. Hence, the spin waves are composed of a dispersive mode and a relatively flat mode, in analogy with the essential character of the modes in Refs. 18, 28. The localized spin becomes weakly dispersive as shown in Fig. 6 with $s_{a}=1$, $S_{c}=6$, $J=1$, $J^{\prime}=0.1$, $D_{a}$=0, and for $D_{c}=0.25$ (green), 0.35 (blue), and 0.45 (red). The energy minimum of the nearly flat mode is 2.0, 2.6, 3.2 meV for $D_{c}$=0.25, 0.35, 0.45, respectively. When the energy minimum of the flat mode is associated with the lowest CFE, the change of $D_{c}$ can be interpreted as that of an applied field $B$ with $D_{c}$ decreasing as the field increases, as seen in Fig. 5, in which $D_{c}=$ 0.45, 0.35, 0.25 correspond to $H\sim$ 0, 5, and 10 T, respectively.
Figure 6: Dispersion relation of the spin wave excitation of Eq. (7) with $s_{a}=1$, $S_{c}=6$, $J=1$, $J^{\prime}=0.1$, $D_{a}$=0. Depending on $D_{c}$, the flat mode changes as red ($D_{c}$=0.45), blue ($D_{c}$=0.35) and green ($D_{c}$=0.45). In comparison, the dispersive mode changes much more weakly with $D_{c}$ and the variation is not visible on this scale. The dispersive mode seems to be black because three cures overlap each other.
@@
While this could be corrected with fuller consideration of the magnetic structure, it may be just as important to go beyond the approximation of single non-interacting magnons. Although we used a local spin injection at the REIG-metal interface in Eq. (17), it is known that the spin Seebeck effect depends on the thickness of the insulator layer [30, 31]. The observed thickness dependence can be attributed to a finite magnon propagation length [30, 31, 32, 33, 34]. The right-hand side of Eq. (17) is proportional to the magnon population from the local spin injection at the interface. When there is a diffusive propagation of magnons in a ferromagnet, the magnon population will acquire an additional population due to the diffusive contribution from the bulk. Moreover, the additional magnon population depends on the damping constant in the ferromagnet [30] measured—the smaller the damping constant, the greater the magnon accumulation. The model in Eq. (7) has two contrasting spin waves, i.e., the almost flat mode and a more dispersive mode. The flat mode corresponds to excitations of the anisotropic rare-earth moments, and as we have seen, its energies are sensitive to the local environment of each site and therefore to any local distortions. It has a small dispersion primarily via exchange through the Fe moments. It therefore has a small group velocity and its damping is most likely large. The dispersive mode has greatest weight on the Fe moments, has a large group velocity and its damping may be rather small. We argue this by similarity of the dispersive mode to the acoustic magnons in YIG, which have extremely small damping. Of course, unlike the acoustic magnon in YIG it develops a gap through hybridization with excitations of the rare-earth, but, as seen in Fig.  6, its energy is much less influenced by fluctuations in the local anisotropy than the flat mode. The diffusion constant $D$ is estimated as $D\propto v_{g}^{2}/\alpha$, with the group velocity of the spin wave $v_{g}$ and its damping constant $\alpha$. Therefore, the dispersive mode will make a relatively large contribution from diffusion in the bulk, whereas the flat mode will contribute very little from diffusion effects. If the geometry allows an additional contribution from diffusion effects, the curves shown in Fig. 7 will be displaced in the negative direction. Under these conditions, the sign change at low temperatures and small fields would not be observable. This may reconcile our predictions with the experimental observation in Ref. 13.
@@ V Summary and Discussions
Incorporating Tb3Fe5O12 (TbIG) into a spin Seebeck effect (SSE) device for thermoelectric generation, the SSE voltage measured in the metal goes to zero at low temperatures. This has been ascribed to the strong single-ion anisotropies of the Tb ions responsible for the opening of a gap in the magnon spectrum [18]. When a magnetic field is applied, one might, by comparison with experiments with other materials, expect the SSE to remain zero as the magnon gap should further increase. Yet for TbIG the SSE instead increases to positive values [13, 14]. We provide an explanation for this based on the effects of crystalline electric fields. The crystal-field excitations (CFE) in TbIG are calculated with a mean field approximation for the effects of magnetic iron moments, using the full set of CF parameters for Tb in Tb3Ga5O12 (TGG) experimentally estimated via neutron scattering in Ref. 21.
@@
In contrast to TGG, where symmetry and the weak interactions between rare-earths allow all the spectra to be the same on each rare-earth $c$ site, in TbIG we must classify the Tb sites into two groups ($\mathbf{C}$ or $\mathbf{C^{\prime}}$ sites). Each site belongs to a set of three, related by 120 degree rotations around the [111] axis for both local axes and the direction of the net magnetic moments of iron with respect to the local axes. Within each group, the crystal-field levels are identical, but the two groups have different canting angles and different energy levels. A given measured crystal-field level can be identified as primarily on either the $\mathbf{C}$ or $\mathbf{C^{\prime}}$ sites, even taking into account weak dispersive effects.
@@
For one group of Tb sites, the $\mathbf{C}$, the lowest CFE decreases when a magnetic field is applied in the [111] crystallographic direction for reasonable values of the mean magnetic field, and Stevens parameters taken from Terbium Gallium Garnet (where the $a$ and $d$ sites are non-magnetic). This can be understood as an extrapolation, after avoided crossings, of the splitting of excited “pseudo-doublets” of TGG in a small external field [21]. As a consequence of this magnetic field dependence of the lowest CFE, the SSE in TbIG can then be enhanced by a magnetic field at low temperatures.
@@
These results, and results reported in Ref. 16, prompted our interest in investigating the expected behaviour of the CFE when the external field is along the crystallographic [100], while this is considered sufficiently weak that the molecular field is still on the [111] direction. Though the clear distinction between the behaviour of levels at the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ sites no longer holds—since the net field lies along a crystallographic axis different from the [111]—our results (details can be found in Appendix B) again show that of the six different $c$-sublattices, three (now sites 1,2 and 5) would feature the slightly unconventional decrease of the lowest CFE with respect to the strength of the field, whilst the remaining three (sites 3,4 and 6) increase with magnetic field strength. This allows us to make predictions for the low-temperature regime where the spin-Seebeck sign change is not yet measured in the presence of a [100] applied field [16].
@@ Acknowledgment
We are indebted to Profs. M. Fujita, T. Kikkawa, and Drs. D. Mannix, S. Geprägs, and J. Thomas-Hunt for many valuable discussions. In particular, we are grateful to Stephan Geprägs and Danny Mannix for sharing unpublished data for transport, magnetic structure and spectroscopy and explaining the difficulties of their interpretation, and for introducing two of us (B.T. and T.Z.) to the subtleties of the subject. This work was supported by JSPS Grant Nos. JP20K03810, JP21H04987, JP23K03291 and the inter-university cooperative research program (No. 202312-CNKXX-0016) of the Center of Neutron Science for Advanced Materials, Institute for Materials Research of Tohoku University. We are grateful to the Global Institute for Materials Research Tohoku that supported the collaboration that led to this work, as did the Reimei program of the Japan Atomic Energy Agency. A part of the computations were performed on supercomputers at the Japan Atomic Energy Agency.
Appendix A Local coordinate systems of the rare-earth sites
@@
The local coordinates $\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{z}_{i}$ are parallel to the 2-fold axes of the $D_{2}$ point group of each $c$-site. Using $\mathbf{x}_{i}=\mathsf{R}_{i}\mathbf{X},\mathbf{y}_{i}=\mathsf{R}_{i}\mathbf{Y},\mathbf{z}_{i}=\mathsf{R}_{i}\mathbf{Z}$, where $\mathsf{R}_{i}$ are the Euler rotation matrices with angles $(\alpha_{i},\beta_{i},\gamma_{i})$ in Table A, the local coordinate systems can be obtained from the cubic axes $\mathbf{X,Y,Z}$, i.e., the ‘laboratory’ coordinate system. State vectors and operators of the quantised angular momenta of the rare-earth moments, are subject to the corresponding Wigner rotation matrices, ${\cal D}(\alpha_{i},\beta_{i},\gamma_{i})=\exp(i\gamma_{i}J_{z})\exp(i\beta_{i}J_{y})\exp(i\alpha_{i}J_{z})$, using the same $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ in Table A.
@@ 6 $\pi/2$ $\pi/4$ $23\pi/$ 2
Table 2: Euler angles $(\alpha_{i},\beta_{i},\gamma_{i})$ for the $c$-sites, as from Ref. 20.
Appendix B Magnetic field dependence of the crystal-field excitations
The magnetic field dependence of the six lowest CF excitations of the single-ion Hamiltonian $H_{CF}-g_{J}\mu_{\mathrm{B}}\,\hat{\mathbf{J}}\cdot\mathbf{B}$, is plotted for “TGG” (not TbIG) in Fig. 8, using $H_{CF}$ from Eq. (2) with parameters from Ref. 21, and $\mathbf{B}$ along the crystallographic [111] direction [20]. At first glance, a contrast emerges between the behaviour of the energies for the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ sites. This grouping is due to the effective projection of the [111] crystallographic axis on the local coordinate systems of the RE $c$-sites, [1,0,$\sqrt{2}$] for the coordinates (21a)-(21c), and [1,$\sqrt{2}$,0] for the coordinates (22a)-(22c[18, 35].

### v3

@@ IV Spin current generation in TbIG at low temperatures
To address under which conditions the CFE could contribute to generating a spin current, we adopt, then, a simplified Hamiltonian that includes the essential low energy behaviour of a relatively localized crystal field level coupled to a single dispersive mode of opposite chirality.
@@
This could be considered as an extension to the model adopted in Ref. 13 (see the supplemental material therein), to include the magnetic field dependence of the CF excitations, with the parameter $D_{c}=D_{c}(B)$ a parametrization of the curve in Fig. (5). There are only two sub-lattices represented in this model, one for the Fe spins $\vec{s}_{a}$ and one for Tb spins $\vec{S}_{c}$. Here the suffixes $a$ and $c$ are to distinguish the distinct Fe and Tb sublattices, respectively, as reminiscent from the Wyckoff positions in garnet crystals. The Fe-Fe and Tb-Fe exchange constants are $J$ and $J^{\prime}$, respectively. While, as we have mentioned, we take values of $D_{c}$ from Fig. 5, which includes effects of the full crystal field potential, we can also consider the model intuitively as having spins are subjected to simplified single-ion anisotropy potentials ($D_{a},D_{c}$) incorporating the axial constraints known from the garnet materials. In practice we neglect the iron anisotropy as being $D_{a}$ much smaller.
@@
where the $\eta$ summation is taken over the $z_{a}=12$ nearest neighbor $a$ sites of one $a$ site, and the $\tilde{\eta}$ summation is of over the $z_{c}=6$ nearest neighbouring $c$ rare-earth sites of one $a$ site. Note that in Eq. (7) the RE-sites interact only with the n.n. Fe-sites and have no interaction within the same sublattice. Hence, the spin waves are composed of a dispersive mode and a relatively flat mode, in analogy with the essential character of the modes in Refs. 21, 31. The localized spin becomes weakly dispersive as shown in Fig. 6 with $s_{a}=1$, $S_{c}=6$, $J=1$, $J^{\prime}=0.1$, $D_{a}$=0, and for $D_{c}=0.25$ (green), 0.35 (blue), and 0.45 (red). The energy minimum of the nearly flat mode is 2.0, 2.6, 3.2 meV for $D_{c}$=0.25, 0.35, 0.45, respectively. When the energy minimum of the flat mode is associated with the lowest CFE, the change of $D_{c}$ can be interpreted as that of an applied field $B$ with $D_{c}$ decreasing as the field increases, as seen in Fig. 5, in which $D_{c}=$ 0.45, 0.35, 0.25 correspond to $H\sim$ 0, 5, and 10 T, respectively.
Figure 6: Dispersion relation of the spin wave excitation of Eq. (7) with $s_{a}=1$, $S_{c}=6$, $J=1$, $J^{\prime}=0.1$, $D_{a}$=0. Depending on $D_{c}$, the flat mode changes as red ($D_{c}$=0.45), blue ($D_{c}$=0.35) and green ($D_{c}$=0.45). In comparison, the dispersive mode changes much more weakly with $D_{c}$ and the variation is not visible on this scale. The dispersive mode seems to be black because three curves overlap each other.
@@
While this could be corrected with fuller consideration of the magnetic structure, it may be just as important to go beyond the approximation of single non-interacting magnons. Although we used a local spin injection at the REIG-metal interface in Eq. (17), it is known that the spin Seebeck effect depends on the thickness of the insulator layer [33, 34]. The observed thickness dependence can be attributed to a finite magnon propagation length [33, 34, 35, 36, 37]. The right-hand side of Eq. (17) is proportional to the magnon population from the local spin injection at the interface. When there is a diffusive propagation of magnons in a ferromagnet, the magnon population will acquire an additional population due to the diffusive contribution from the bulk. Moreover, the additional magnon population depends on the damping constant in the ferromagnet [33] measured—the smaller the damping constant, the greater the magnon accumulation. The model in Eq. (7) has two contrasting spin waves, i.e., the almost flat mode and a more dispersive mode. The flat mode corresponds to excitations of the anisotropic rare-earth moments, and as we have seen, its energies are sensitive to the local environment of each site and therefore to any local distortions. It has a small dispersion primarily via exchange through the Fe moments. It therefore has a small group velocity and its damping is most likely large. The dispersive mode has greatest weight on the Fe moments, has a large group velocity and its damping may be rather small. We argue this by similarity of the dispersive mode to the acoustic magnons in YIG, which have extremely small damping. Of course, unlike the acoustic magnon in YIG it develops a gap through hybridization with excitations of the rare-earth, but, as seen in Fig.  6, its energy is much less influenced by fluctuations in the local anisotropy than the flat mode. The diffusion constant $D$ is estimated as $D\propto v_{g}^{2}/\alpha$, with the group velocity of the spin wave $v_{g}$ and its damping constant $\alpha$. Therefore, the dispersive mode will make a relatively large contribution from diffusion in the bulk, whereas the flat mode will contribute very little from diffusion effects. If the geometry allows an additional contribution from diffusion effects, the curves shown in Fig. 7 will be displaced in the negative direction. Under these conditions, the sign change at low temperatures and small fields would not be observable. This may reconcile our predictions with the experimental observation in Ref. 16.
@@ V Summary and Discussions
Incorporating Tb3Fe5O12 (TbIG) into a spin Seebeck effect (SSE) device for thermoelectric generation, the SSE voltage measured in the metal goes to zero at low temperatures. This has been ascribed to the strong single-ion anisotropies of the Tb ions responsible for the opening of a gap in the magnon spectrum [21]. When a magnetic field is applied, one might, by comparison with experiments with other materials, expect the SSE to remain zero as the magnon gap should further increase. Yet for TbIG the SSE instead increases to positive values [16, 17]. We provide an explanation for this based on the effects of crystalline electric fields. The crystal-field excitations (CFE) in TbIG are calculated with a mean field approximation for the effects of magnetic iron moments, using the full set of CF parameters for Tb in Tb3Ga5O12 (TGG) experimentally estimated via neutron scattering in Ref. 24.
@@
In contrast to TGG, where symmetry and the weak interactions between rare-earths allow all the spectra to be the same on each rare-earth $c$ site, in TbIG we must classify the Tb sites into two groups ($\mathbf{C}$ or $\mathbf{C^{\prime}}$ sites) determined by the direction of net magnetization, assumed parallel to [111]. Each site belongs to a set of three, related by 120 degree rotations around the [111] axis for both local axes and the direction of the net magnetic moments of iron with respect to the local axes. Within each group, the crystal-field levels are identical, but the two groups have different canting angles and different energy levels. A given measured crystal-field level can be identified as primarily on either the $\mathbf{C}$ or $\mathbf{C^{\prime}}$ sites, even taking into account weak dispersive effects.
@@
For one group of Tb sites, the $\mathbf{C}$, the lowest CFE decreases when a magnetic field is applied in the [111] crystallographic direction for reasonable values of the mean magnetic field, and Stevens parameters taken from Terbium Gallium Garnet (where the $a$ and $d$ sites are non-magnetic). This can be understood as an extrapolation, after avoided crossings, of the splitting of excited “pseudo-doublets” of TGG in a small external field [24]. To see this, in Fig. 9 of Appendix C, note that the decreasing part of the lowest excitation level (red, on-line) can be seen to be the approximate continuation of the second level (blue, on-line) after an avoided crossing at lower fields. As a consequence of this magnetic field dependence of the lowest crystal field level, the Spin Seebeck Effect in TbIG can then be enhanced by a magnetic field at low temperatures.
@@
Results reported in Ref. 19, prompted our interest in further investigating the expected behaviour of the CFE when the external field is along the crystallographic [100], while this is considered sufficiently weak that the molecular field is still on the [111] direction. Though the clear distinction between the behaviour of levels at the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ sites no longer holds—since the net field lies along a crystallographic axis different from the [111]—our results (details can be found in Appendix C) show that of the three different $\mathbf{C}$ sublattices contributing to the lowest crystal field level, two (sites 1,2) still feature the slightly unconventional decrease of the lowest CFE with respect to the strength of the field, but site 3 now increases. For the $\mathbf{C^{\prime}}$, 4 and 5 also decrease with external magnetic field while 6 increases. As the energies of the $\mathbf{C^{\prime}}$ are higher even in zero external field, this will probably not determine the spin-Seebeck effect at low temperatures. From the decrease of the energies of the two $\mathbf{C}$ sites, this allows us to predict that at low temperatures, the sign change in the spin-Seebeck effect should also appear in the presence of a [100] applied field, but more weakly than for the [111]. While the spin-Seebeck effect has already been measured with the field in this direction [19], results have not yet been reported at low temperatures.
@@ Acknowledgment
We are indebted to Profs. M. Fujita, T. Kikkawa, and Drs. D. Mannix, S. Geprägs, and J. Thomas-Hunt for many valuable discussions. In particular, we are grateful to Stephan Geprägs and Danny Mannix for sharing unpublished data for transport, magnetic structure and spectroscopy and explaining the difficulties of their interpretation, and for introducing two of us (B.T. and T.Z.) to the subtleties of the subject. This work was supported by JSPS Grant Nos. JP20K03810, JP21H04987, JP23K03291 and the inter-university cooperative research program (No. 202312-CNKXX-0016) of the Center of Neutron Science for Advanced Materials, Institute for Materials Research of Tohoku University. B.T. acknowledges partial support from the project PRIN 2022 - 2022XK5CPX (PE3) SoS-QuBa - ‘Solid State Quantum Batteries: Characterization and Optimization’. We are grateful to the Global Institute for Materials Research, Tohoku University that supported the collaboration that led to this work, as did the Reimei program of the Japan Atomic Energy Agency. A part of the computations were performed on supercomputers at the Japan Atomic Energy Agency. In the mean field calculation, we used a Python class file of the Wigner rotation included in PyCrystalField [40]. The single-ion results used to benchmark the theory and calculations in this work were obtained utilising an extended version of a custom code initially developed by B.T. for Refs. [24, 21].
Appendix A Position of Tb site in a cubic unit cell
To clarify the difference between the $\mathbf{C}$ (with $i=1,2,3$) and the $\mathbf{C^{\prime}}$ (with $i=4,5,6$) sites, the Tb sites in a cubic unit cell are shown below. The atomic coordinate of Tb sites are listed in Ref. 29 as,
{tblr}
colspec = —X[2,c]—X[1,c]—X[1,c]—X[1,c]—, width = 1.0 $\mathbf{C}$ sites x y z
$i=1$ 0 1/4 1/8
0 3/4 7/8
1/2 3/4 5/8
1/2 1/4 3/8
$i=2$ 1/4 1/8 0
3/4 7/8 0
3/4 5/8 1/2
1/4 3/8 1/2
$i=3$ 1/8 0 1/4
7/8 0 3/4
5/8 1/2 3/4
3/8 1/2 1/4
{tblr}
colspec = —X[2,c]—X[1,c]—X[1,c]—X[1,c]—, width = 01.0 $\mathbf{C^{\prime}}$ sites x y z
$i=4$ 0 3/4 3/8
0 1/4 5/8
1/2 1/4 7/8
1/2 3/4 1/8
$i=5$ 3/4 3/8 0
1/4 5/8 0
1/4 7/8 1/2
3/4 1/8 1/2
$i=6$ 3/8 0 3/4
5/8 0 1/4
7/8 1/2 1/4
1/8 1/2 3/4
Table 2: The atomic coordinates of the Tb $c$-sites, as from Ref. 29. The $i$ labels are attributed to the $c$ positions according to their associated local coordinate system, here in Eqs.21-22.
These sites are plotted in Fig. 8, where the iron and oxygen sites are omitted for clarity. We note that if we chose another, but equivalent direction for the net magnetization, e.g. $[11\overline{1}]$, the groupings of sites into perpendicular triangles and $\mathbf{C}$ and $\mathbf{C^{\prime}}$ classes would be different.
Figure 8: The Tb site in a unit cell from a view point in the [111] (panel a) and [100] (panel b) directions. The numbers correspond to the index $i=1\sim 6$ of the $c$-sites. The 1, 2, 3 sites form triangles in [111] planes, as do the 4, 5, 6, and this determines classification as $\mathbf{C}$ and $\mathbf{C^{\prime}}$. Figures are made using VESTA software [39].
Appendix B Local coordinate systems of the rare-earth sites
@@
The local coordinates $\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{z}_{i}$ are parallel to the 2-fold axes of the $D_{2}$ point group of each $c$-site. Using $\mathbf{x}_{i}=\mathsf{R}_{i}\mathbf{X},\mathbf{y}_{i}=\mathsf{R}_{i}\mathbf{Y},\mathbf{z}_{i}=\mathsf{R}_{i}\mathbf{Z}$, where $\mathsf{R}_{i}$ are the Euler rotation matrices with angles $(\alpha_{i},\beta_{i},\gamma_{i})$ in Table B, the local coordinate systems can be obtained from the cubic axes $\mathbf{X,Y,Z}$, i.e., the ‘laboratory’ coordinate system. State vectors and operators of the quantised angular momenta of the rare-earth moments, are subject to the corresponding Wigner rotation matrices, ${\cal D}(\alpha_{i},\beta_{i},\gamma_{i})=\exp(i\gamma_{i}J_{z})\exp(i\beta_{i}J_{y})\exp(i\alpha_{i}J_{z})$, using the same $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ in Table B.
@@ 6 $\pi/2$ $\pi/4$ $23\pi/$ 2
Table 3: Euler angles $(\alpha_{i},\beta_{i},\gamma_{i})$ for the $c$-sites, as from Ref. 23.
Appendix C Magnetic field dependence of the crystal-field excitations
The magnetic field dependence of the six lowest CF excitations of the single-ion Hamiltonian $H_{CF}-g_{J}\mu_{\mathrm{B}}\,\hat{\mathbf{J}}\cdot\mathbf{B}$, is plotted for “TGG” (not TbIG) in Fig. 9, using $H_{CF}$ from Eq. (2) with parameters from Ref. 24, and $\mathbf{B}$ along the crystallographic [111] direction $\mathbf{B}\,[111]$ [23]. Below, boldface denotes a vector and non-boldface is the magnitude of a vector. At first glance, a contrast emerges between the behaviour of the energies for the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ sites. This grouping is due to the effective projection of the [111] crystallographic axis on the local coordinate systems of the RE $c$-sites, [1,0,$\sqrt{2}$] for the coordinates (21a)-(21c), and [1,$\sqrt{2}$,0] for the coordinates (22a)-(22c[21, 38].